2.50 · 104 moving at a speed of 4.00 m / s
kg wheel rail
were moving in the same direction together at an initial speed of 2.00 m / s, each of which was 2.50 · 104
kg
collided with a series of three wheelchairs interlocking with the mass
The four-wheeler began to move in one direction. A) Post-conflict 4
What is the speed of the wheelchair? B) How much mechanical energy was lost in the collision?
is it
1
Expert's answer
2020-05-04T12:46:53-0400

a)


"mv+3mu=(4m)V"

"V=0.25(v+3u)=0.25(4+3(2))=2.5\frac{m}{s}"

b)


"\Delta E=0.5mv^2+0.5(3mu^2)-0.5(4mV^2)\\=0.5m(v^2+3u^2-4V^2)"

"\Delta E=0.5(25000)(4^2+3(2)^2-4(2.5)^2)=37500\ J"